# floating point operation IVP initial value problem

### v1

@@ VI-D Near Optimal Solutions
Furthermore, the optimal number of samples also depends on the annealing time $T_{A}$. As a preliminary step, we investigate the distribution of the relative error $L$ for the best embedding and chain strength combinations found in the previous section for different $T_{A}$, cf. Fig. 5a. We find that the mean and the quantiles of the distributions of $L$ are, as expected, lower for $T_{A}=50\mu s$ than for $T_{A}=1\mu s$. Hence, the optimal $T_{A}$ for the whole loop can not be assessed by just looking at the $TTS$ for the optimization problem alone. The distributions of $L$ also reveal that the variance in $L$ becomes smaller for larger grids. More surprisingly, the mean of the distributions for case 57 is lower than for case 30. Hence, we can hope that sampling partitionings on current (or near-term) QPUs for larger grids will also return near-optimal solutions that are good enough to provide simulation speed-ups in competitive time.
Near-optimal solutions might be less intuitive than optimally partitioned sub-networks. For case 14, first, the cut becomes larger with increasing $L$; however, the components and sub-network sizes remain (roughly) balanced, cf. Fig. 5b. The increase in cut size leads to disconnected sub-networks.
@@ VII Conclusion and Outlook
VIII Acknowledgements
The paper was written as part of the project ”Quantum-based Energy Grids (QuGrids)”, which is receiving funding from the programme ”Profilbildung 2022”, an initiative of the Ministry of Culture and Science of the State of North Rhine-Westphalia. The authors gratefully acknowledge the Jülich Supercomputing Centre (https://www.fz-juelich.de/en/ias/jsc/systems/supercomputers/apply-for-computing-time/juniq) for funding this project by providing computing time on the D-Wave Advantage™ System JUPSI through the Jülich UNified Infrastructure for Quantum computing (JUNIQ), which has received funding from the German Federal Ministry of Education and Research (BMBF) and the Ministry of Culture and Science of the State of North Rhine-Westphalia. The authors also thank Manuel Dahmen for the fruitful discussions. The sole responsibility for the content of this publication lies with the authors.

### v2

@@ VI-D Near Optimal Solutions
Furthermore, the optimal number of samples also depends on the annealing time $T_{A}$. As a preliminary step, we investigate the distribution of the relative error $L$ for the best embedding and chain strength combinations found in the previous section for different $T_{A}$, cf. Fig. 5a. We find that the mean and the quantiles of the distributions of $L$ are, as expected, lower for $T_{A}=50\mu s$ than for $T_{A}=1\mu s$. Hence, the optimal $T_{A}$ for the whole loop can not be assessed by just looking at the $TTS$ for the optimization problem alone. Interestingly, the mean and variance of the distributions of $L$ for case 57 are lower than those for case 30, although solving case 57 up to optimality is harder than solving case 30 for QA. We assume that this alleged paradox can be resolved by comparing the topologies and generator locations of case 30 and case 57, as shown in Fig. 2. In case 57, the generators are all concentrated in the “southwest” of the grid. As a result, any partition that seeks to balance both the components, particularly the costly generators and the sizes of the sub-networks, must cut diagonally from the southwest to the northeast. Furthermore, the northeast region consists of only a few large cycles, further limiting “good” partition choices. This translates to a significant gap between the energy (objective value) levels of $Q_{part.}$ corresponding to “good” cuts along the southwest-to-northeast axis and the energy levels for all other “bad” cuts, where the generators would be strongly imbalanced. Therefore, during the anneal, the system is unlikely to get excited to the higher energy levels associated with “bad” cuts along any other axis. On the other hand, case 30 is more homogeneous, and thus we expect that energy levels are more evenly distributed. Consequently, transitions or cascades of transitions to higher energy levels corresponding to “bad” partitionings are more likely.
In general, we can hope that sampling partitionings on current (or near-term) QPUs for larger grids will also return near-optimal solutions that are good enough to provide simulation speed-ups in competitive time. However, near-optimal solutions might be less intuitive than optimally partitioned sub-networks. For case 14, first, the cut becomes larger with increasing $L$; however, the components and sub-network sizes remain (roughly) balanced, cf. Fig. 5b. The increase in cut size leads to disconnected sub-networks.
VI-E Rescaling the QUBO to improve performance
The three QUBO objective functions of $Q_{part}$ have been designed to estimate the overhead or (squared) idle time associated with each of the three steps in every iteration of the parallel simulation. So far, $Q_{part}$ is a simple sum of the three objectives, effectively balancing the overheads and squared idle times, see Eq. 18. However, in the context of linear scalarization (cf. Eq. 12), the individual objectives can be rescaled by introducing scalars as
$Q_{part}({\boldsymbol{\lambda}})=\lambda_{comp.}Q_{comp.}+\lambda_{cut}Q_{cut}+\lambda_{net.}Q_{net.}.$
Rescaling changes the importance of the single objectives and altering the energy landscape of the problem and thus eventually the optimal partitioning. On the other hand, any reformulation, and thus any rescaling of the single objectives, can also influence the solution time of solvers involving heuristics. For QA rescaling can modify the separation of the energy levels and thus transition probabilities to excited states.
For case 30, we observe that increasing the importance of $Q_{cut}$, in particular setting $\lambda_{cut}=2$ while keeping $\lambda_{comp.\backslash net.}=1$, significantly improves the performance of QA, see Fig. 6 a, while the optimal partitioning is not affected, see subfigure b. For $\lambda_{cut}=2$ the optimal solution is found for all tested chain strengths. The fastest $TTS$ is $0.21s$, which constitutes an improvement by a factor of 110 compared to the fastest $TTS$ with $\lambda_{cut}=1$. Further research is needed to test the effect of rescaling across different test cases, as well as, the trade-off between performance and optimal partitionings.
Figure 6: Effect of rescaling the energy of $Q_{cut}$ for case 30. a: Comparison of the performance ($TTS$) and solution quality ($L_{min}$) obtained by QA between $\lambda_{cut}=1$ and $\lambda_{cut}=2$. For each $RCS$ value only the minimal $L_{min}$ with the fastest $TTS$ among all 10 embeddings is shown here. b: Optimal partitioning for $\lambda_{cut}=1$ and $\lambda_{cut}=2$.
@@ VII Conclusion and Outlook
VIII Code Availability
The data and code that support the findings of this article are openly available [66].
IX Acknowledgements
The paper was written as part of the project “Quantum-based Energy Grids (QuGrids)”, which is receiving funding from the programme “Profilbildung 2022”, an initiative of the Ministry of Culture and Science of the State of North Rhine-Westphalia. The authors gratefully acknowledge the Jülich Supercomputing Centre (https://www.fz-juelich.de/en/ias/jsc/systems/supercomputers/apply-for-computing-time/juniq) for funding this project by providing computing time on the D-Wave Advantage™ System JUPSI through the Jülich UNified Infrastructure for Quantum computing (JUNIQ), which has received funding from the German Federal Ministry of Education and Research (BMBF) and the Ministry of Culture and Science of the State of North Rhine-Westphalia. The authors also thank Manuel Dahmen for the fruitful discussions. The sole responsibility for the content of this publication lies with the authors.
X Appendix
Figure 7: Solution quality across different embedding and chain strength combinations for cases 14, 30, and 57. The data has been “projected” into a 2d plane for Fig. 4 a.